Solve 2x^2 − 12x + 4 = 0 by completing the square, expressing the result in simplest radical form.
Answer by mananth(16946)

2x^2-12x+4 =0
x^2-6x+2=0
x^2-6x+9 -7=0
(x-3)^2 - (sqrt7)^2
= (x-3+sqrt7)(x-3-sqrt7)